OpenStudy (anonymous):

Solve the equation: 9x-2=4(x+3)

OpenStudy (saifoo.khan):

9x - 2 = 4x + 12

OpenStudy (saifoo.khan):

5x = 14 x = 14/5

OpenStudy (anonymous):

9x - 2 = 4(x + 3) 9x - 2 = 4x + 12 5x = 14 x = 5/14

OpenStudy (saifoo.khan):

hick, u did a mistake.. 14/5

OpenStudy (anonymous):

yeah, i'm obviously wrong. 14/5. :-)

OpenStudy (saifoo.khan):

lol, np!

OpenStudy (llort):

x=14/5

OpenStudy (llort):

or 5x-14=0